-If Suzanne weighs 62 kg what is the testing weight for a Wingate test?

-What is Suzanne’s peak power on her Wingate test if she achieved 11 revolutions in her first 5 second interval, and the meters traveled are 6 m.

-When performing a one repetition max test what is the first load that should be added to the bar?
Answer #1

Testing weight for Wingate test = 7.5 % of participant's weight

= 7.5 /100 * 62 = 4.65 kg

Peak power= Force * Distance travelled / Time(in minutes)

= 4.65 * (6*11) * 12/6.12

= 601.8 W

First loss to be added = 4.65 kg